OpenStudy (anonymous):

The integral of 1/(x^2+9)^(3/2)dx Help!

OpenStudy (anonymous):

Yes, but this problem deals with the the In of a function...

OpenStudy (anonymous):

the last one wasn't correct, not trolling. Just trying to help. i reposted on your old question

OpenStudy (anonymous):

What do you mean it wasnt correct? I said tan^5x/5 +c

OpenStudy (anonymous):

your post said tan^2x/2+c, looks like just a typo

OpenStudy (anonymous):

Lol yeah it must have been a typo :P Sorry and thank you :)

OpenStudy (anonymous):

So do you have any advice for this problem?

OpenStudy (anonymous):

sorry i've got no luck with this one yet...>.<

OpenStudy (anonymous):

Okay...Have you tried Anti-log?

OpenStudy (anonymous):

you mean like 1/x =ln(x)*x'?

OpenStudy (anonymous):

Yeah....Its the first one. :)

OpenStudy (anonymous):

wolfram got the answer of \[\frac{ x }{ 9\sqrt{x^2+9}}\]...i havent been able to get that

OpenStudy (anonymous):

Yeah I saw that on my calculator but I havent been able to get to that answer either :/

OpenStudy (anonymous):

Okay nvm dude...I solved it :)

OpenStudy (anonymous):

nice, what'd you use?

OpenStudy (anonymous):

I set x=Squ(9)sectheta and dx=Squ(9)(sectheta)(tantheta)(dtheta)

OpenStudy (anonymous):

I used Squ(x^2-a^2) which means x=(a)sectheta

OpenStudy (anonymous):

And yeah :P I did a lot of work after that. But thats how I started and I finished :)

OpenStudy (anonymous):

ahhh nice

OpenStudy (anonymous):

Yep :) Thanks for the help on the last problem :P